Write a quadratic model for (-4,8) (-1,5) (1,13)

Found 2 solutions by ankor@dixie-net.com, Alan3354:
Answer by ankor@dixie-net.com(22740)

Write a quadratic model for (-4,8) (-1,5) (1,13)
Using the form ax^2 + bx + c = y, write an equation for each x,y pair
(-4,8)
(-4^2)a - 4x + c = 8
16a - 4b + c = 8
:
(-1,5)
a - b + c = 5
:
(1,13)
a + b + c = 13
:
Using elimination on the last two equations
a + b + c = 13
a - b + c = 5
--------------subtraction eliminates a & c, find b
0 + 2b + 0 = 8
2b = 8
b = 4
:
Replace b in the 1st and 3rd equations
16a -4(4) + c = 8
16a + c = 8 + 16
16a + c = 24
:
a + 4 + c = 13
a + c = 9
Use elimination on these two equations
16a + c = 24
a + c = 9
------------subtraction eliminates c
15a = 15
a = 1
Find c using a + c = 9
1 + c = 9
c = 8
the equation: Y = x^2 + 4x + 8

Answer by Alan3354(69443)

Write a quadratic model for (-4,8) (-1,5) (1,13)
========
There are 2 possible parabolas for these 3 points.
---
y = ax^2 + bx + c is one of them, and a solution is given.
The other is x = ay^2 + by + c
----
Use the same method to find it.